OpenStudy (anonymous):

What is the y-intercept for the graph of the equation 4x - 4y = 9?

mathslover (mathslover):

take x = 0 . 4(0)-4y=9 -4y=9 y=\(\frac{-9}{4}\)

mathslover (mathslover):

hence y intercept is (0,-9/4)

OpenStudy (goformit100):

@mathslover is correct